An open flask sitting in a lab refrigerator looks empty, but is actually filled with a mixture or gases called air. If the flask volume is 1.00 L, and the air is at standard temperature and pressure, how many gaseous molecules does the flask contain?


1
Expert's answer
2021-02-05T04:04:26-0500

n= N/NA, N= nxNA.

n= V/22.4 = 0,0446428571 moles.

N= 0,0446428571 x 6.02 x 10^23 = 2,6875e22 .